Grade 12th passGeneral Physics

A body of mass 5kg is acted upon by two perpandicular forces 16N and 12N. Find magnitude and direction of force and also find accelration??.

Harpreet
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
12 N5kg is acted upon by two perpandicular forces 16N and 
 
Fnet = root(16^2 + 12^2) = 20 Newton
cos@ = 16/20 = 4/5 
@ = 36.8 degree 
acceleration = 2.4i + 2.2j